A stone is thrown straight upward with a velocity of 30 m/s. What was its maximum height?, b) What is the time taken to reach the maximum height?, c) What is the velocity of the stone when it is 60 m from the ground? (Assume the ground as the initial frame of reference)

1
Expert's answer
2022-03-18T09:07:55-0400

(a)


"h_{\max}=\frac{v_0^2}{2g}=\frac{30^2}{2\times 9.8}=46\:\rm m"

(b)


"t=\frac{v_0}{g}=\frac{30}{9.8}=3.1\:\rm s"

(c) impossible